The healthcare provider prescribes amoxicillin 125 mg for a child weighing 25 lb. The pharmacy supplies a dosage strength of 250 mg/5mL. Safe (usual) does is 20-40 mg/kg in 3 divided doses. Is this does safe? Please circle one as your answer:

safe, but ineffective

safe

not safe.

If dose above is safe, how many mL's dose would the nurse give?

____________mL

PLEASE PLEASE SHOW WORK IN DIMENSIONAL ANALYSIS ONLY THANK YOU
Answer #1

Prescribed amoxicillin= 125 mg

For the patient weighing 25 lb= 25/2.2= 11.36 kg

Safe recommended dosing= 20-40 mg/kg in 3 divided dose

= 20X 11.36 mg to 40 X 11.36 mg in 3 divided dose

=227.2 mg to 454.4 mg in 3 divided dose

Since the ordered dose i.e 125 mg is within the safe recommended dose range i.e. 227.2 mg to 454.4 mg

Hence it is safe dosing.

Available dose strength= 250 mg/5 ml

=50 mg per ml

Hence for the ordered 125 ml required ml= 125/50 ml

= 2.5 ml